OpenStudy (anonymous):

in how many ways can six people be lined up for a photograph of ralph must stand immediately to the left of bernice

jimthompson5910 (jim_thompson5910):

Hint: If ralph must stand immediately to the left of bernice, then the grouping of ralph and bernice forms one "person" so to speak. This is because ralph must always be to the left of bernice and you can think of these two slots as one big slot so you really have 6-1 = 5 people total and you're ordering them 5 different ways

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

does that make sense?

OpenStudy (anonymous):

wait there's more. can u contitnue?>

jimthompson5910 (jim_thompson5910):

how many ways are there to order 5 people?

OpenStudy (anonymous):

5*4*3*2*1 ok thanks haha

jimthompson5910 (jim_thompson5910):

yep 120 ways

OpenStudy (anonymous):

hey what about this one? in how many ways can nine people be lined up for a photograph if clark and billy must not stand next to each other

jimthompson5910 (jim_thompson5910):

Let's count how many ways clark and billy can stand next to each other So there are 2*8! = 80640 ways for them to do this

jimthompson5910 (jim_thompson5910):

There are 9! = 362880 different ways to order 9 people So there are 9! - 2*8! = 362880 - 80640 = 282240 ways for them to line up where clark and billy must not stand next to each other

OpenStudy (anonymous):

wait so why is it not just 2*8! ?

jimthompson5910 (jim_thompson5910):

because 2*8! is the number of ways to order them where clark and billy are next to each other

jimthompson5910 (jim_thompson5910):

basically you combine clark and billy into one person to go from 9 people to 8 people

OpenStudy (anonymous):

oh I c... so 2*8! is just the way them two can order. but u subtract to also include the other members. ok tahnks!!!

jimthompson5910 (jim_thompson5910):

yep, yw

OpenStudy (anonymous):

oh wait so can u tell me how u know to times 2?